The Character of Gargantua as a Vehicle for Comedy

Marc Horowitz Comedy — 21.021 March 4, 1992

In The Birth and Early Upbringing of Gargantua, Rabelais uses the character of Gargantua as the subject of his humor. He relates the events surrounding Gargantua’s childhood, using them as a vehicle for many common kinds of humor. The general theme of Rabelais’s stories is that of nature gone awry. Gargantua does all the things a normal child would, but he does them bigger, earlier, and faster. The reader laughs, knowing that what is happening is quite impossible.

The most common technique used by Rabelais is to take an everyday situation, and exaggerate it. In the first few stories, which relate Gargantua’s mother’s pregnancy, many events are described which are beyond everyday parameters. Her pregnancy is eleven months long. At lunch the day she gives birth, there is a feast, at which the tripe of 367,014 oxen recently slaughtered are eaten. Gargantua’s mother herself eats 1434 gallons of tripe. Of course, baby Gargantua was no less of an eater. Each day, he required the milk of 17,913 cows. Rabelais does not allow the reader to get used to this world of exaggeration. After mentioning several huge numbers, he moves back into what we would consider reality, and declares that it is “scandalous and offensive to pious ears” (p. 72) to consider that it might be possible for Gargantua’s mother to generate over 58,000 gallons of breast milk. He acheives humor by contradicting realit, and acheives humor again by contradicting his own contradiction. Then, Rabelais continues to contradict reality once more. To make Gargantua clothing, the tailors use several miles of cloth. Gargantua wears a gold chain weighing over 80 tons. Exaggeration is a simple, effective device of which Rabelais makes repeated use in his stories.

Gargantua at various times in the story performs some seemingly physically impossible feats. When he is born, rather than being born through the normal channels, he “entered the hollow vein and climbed over the diaphragm to her shoulders … and there he took the left hand path and came out by the left ear.” (p. 69) At this point, one would expect the baby to start crying. Instead Gargantua cries out, in French (this is a French story), for a drink. Rabelais’s imagery is very good, and we laugh, considering how absurd such actions are.

In describing Gargantua and his childhood, Rabelais finds many opportunities to satirize the establishment, especially literature and religion. In many places, he seeks to add credibility to his story by citing classic writers from Greek literature and the Bible. For instance, after describing the story of Gargantua’s birth, he says that “a man of good sense, ought always to believe what anybody tells him and whatever he believes in print” (p. 69), then goes on to cite the Bible as proof of this (p. 69).

Going through such effort to “prove” a story the author and the reader know is false is humorous. On a deeper level, it causes the reader to laugh at the implication that the Bible itself might contain some fallacies. If the reader still is not convinced, Rabelais cites a series of Greek myths and a book of science to demonstrate that such a birth is possible. In between these literary references, he mentions some other books which are less established, such as his own work, On the Dignity of Codpieces. By juxtaposing this book (whether it is an actual book, or merely a fictional work with an amusing title), Rabelais pokes fun at the literary establishment of which he is a part.

Another target of Rabelais’s satire is the Church’s educational system. The system of that period spent much time reading and rereading books. Lessons were repeated many times. Of course, for Gargantua, more is better. He spends five years learning his alphabet, and eleven years studying grammar and morals. This system is clearly not one of which Rabelais approves. He makes his disapproval clear by ridiculing how much time is wasted on the simplest subjects.

Some of Gargantua’s behavior is not impossible, but farfetched enough that his actions are quite unlikely to happen. At the age of five, Gargantua has already reached sexual maturity and is flirting with his nurses, while they flirt with him. “One of them would call it my little spout, another my peg, another my coral-branch, another my stopper, my plug, my centerbit, my ramrod, my gimlet, my trinket, my rough-and-ready-stiff-and-steady, my prop, my little red sausage, my little booby prize.” (p. 80) These are hardly terms which one would expect nurses to be using about a five-year-old. Rabelais also tells the story of how Gargantua has a serious argument with his father about rump-wiping. Gargantua discusses the relative merits of various implements of rump-wiping, most of which one would never imagine using for such a purpose. These included ladies’ clothing, various animals, herbs, linens, furniture, and other various and sundry items. Of course, he tried paper, of which he says

“He always leaves bait on his balls Who his dirty arse with paper mauls.”

Imagine, using paper to wipe one’s rump. This is another of Rabelais’s satirical “absurd ideas” among truly absurd ideas. Of course, Gargantua concludes that the best rump-wiper of all is a down goose. He further claims that the bliss of the Greek gods was not due to any of the traditional mythological reasons, but to the fact that they wiped their rumps with down gooses. Again, Rabelais mentions that this is “also... the opinion of Master John of Scotland.” Although this reference is meaningless to me, I assume that in Rabelais’s time, this would be an amusing reference to a non-authority on ancient Greece. Again, Rabelais is ridiculing literary classics.

The general pattern of all the comedy in this story is that of taking everyday occurrences, and replacing them or combining them with unusual occurrences. Gargantua is seen as an example of everything a normal person is, but more. In describing what happens to Gargantua, the standard flow of reality is distorted, revealing the humorous aspects of Rabelais’s society.